Eberhart Manufacturing has projected sales of $145 million next year. Costs are expected to be $81 million and net investment is expected to be $15 million. Each of these values is expected to grow at 14 percent the following year, with the growth rate declining by 2 percent per year until the growth rate reaches 6 percent, where it is expected to remain indefinitely. There are
5.5 million shares of stock outstanding and investors require a return of 13 percent return on the company’s stock. The corporate tax rate is 40 percent.

a. What is your estimate of the current stock price?
b. Suppose instead that you estimate the terminal value of the company using a PE
multiple. The industry PE multiple is 11. What is your new estimate of the company’s
stock price?
1
Expert's answer
2018-02-06T07:40:07-0500
Sales = $145 million, Costs = $81 million, net investment = $15 million, g = 14% (declining by 2 percent per year until the growth rate reaches 6 percent, where it is expected to remain indefinitely), 5.5 million shares of stock outstanding, r = 13%, t = 40%.
a. What is your estimate of the current stock price?
P = D0*(1 + g)/(r - g) = (145 - 81 - 15)/5.5*(1 + 0.14 - 0.02)/(0.13 - 0.12) = $997.8.
b. P = PE*EPS = 11*(145 - 81 - 15)/5.5 = $98.